# LAW 2025: Bridging Language, Agent, and World Models

### v3

@@ Abstract
Reinforcement learning based post-training of large language models (LLMs) has recently gained attention, particularly following the release of DeepSeek R1 (guo2025deepseek), which applied GRPO for fine-tuning. Amid the growing hype around improved reasoning abilities attributed to RL post-training, we critically examine the formulation and assumptions underlying these methods. We start by highlighting popular structural assumptions made in modeling LLM training as an MDP, and show how they lead to a degenerate MDP that doesn’t quite need the RL/GRPO apparatus. The two critical structural assumptions include (1) making the MDP states be just a concatenation of the actions with states becoming the context window and the actions becoming the tokens in LLMs and (2) splitting the reward of a state-action trajectory uniformly across the trajectory. Our comprehensive analysis demonstrates that, due to these simplifying assumptions, the standard approach is effectively equivalent to outcome-driven supervised learning. Our experiments on benchmarks including GSM8K and Countdown using Qwen-2.5 base models and Llama-3.2 instruct models show that Filtered Iterative SFT, incorporating both positive and negative samples, achieves performance comparable to GRPO-based training. We also show that these structural assumptions indirectly incentivize RL to generate longer sequences of intermediate tokens which in turn feeds into the narrative of “RL incentivizing thinking because it generates longer thinking traces.” We continue to believe that RL writ large can be a great tool for post-training LLMs, and hope that our analysis of the limiting assumptions in the currently popular RL framework encourages work that goes beyond patching the symptoms of faulty assumptions and towards improving the underlying formulation.
@@ 1 INTRODUCTION
In this work, we critically examine the foundations of applying RL to LLM post-training through the lens of the specific MDP formulation that has been popularly used to model LLMs as agents  (guo2025deepseek). Under this formulation, each state is a sequence of previously generated tokens, transitions are deterministic concatenations of tokens, rewards are assigned only at the terminal state based on external verification of correctness, and the credit assignment problem is short-circuited by effectively dividing this terminal reward across all tokens (or actions) equally. These structural assumptions sharply distinguish the setting from the general MDP framework typically considered in RL, and drive several wrong intuitions.
@@
Our contributions are twofold. Firstly, we theoretically show that under these structural assumptions, the full machinery of RL methods in LLMs post-training is somewhat redundant. By decoupling and simplifying the GRPO objective for trajectories labeled as positive (correct solutions) and negative (incorrect solutions), we demonstrate that iterative supervised fine-tuning (SFT) guided by an external verifier yields comparable performance. This equivalence implies that RL might not be imparting genuinely novel reasoning capabilities to the LLM, but rather leveraging latent capabilities inherent within the pretrained models.
@@
Secondly, we empirically demonstrate that the terminal reward distribution assumption, wherein the relative advantage score is uniformly allocated across all tokens in a response, directly contributes to the increasing response lengths observed throughout training. Contrary to the claims made in the DeepSeek-R1 (guo2025deepseek) paper, which attributes the increased response length at test-time to computational scaling, self-reflection, self-verification, and back-tracking, we show that the primary driver of response elongation is this uniform credit distribution. Since each token is incentivized equally without considering the "return-to-go" from any intermediate state, RL-based methods like GRPO inadvertently encourage verbose outputs.
@@
This paper thus aims to clarify the misconceptions regarding the role of RL in enhancing LLM capabilities by providing a comprehensive theoretical and empirical analysis showing that idiosyncratic assumptions made in formulating the MDP influence model behavior during post training. A key implication of our analysis is that misunderstanding the consequences of these structural assumptions has led to research directions that attempt to fix superficial symptoms such as introducing length penalties (arora2025training) or selectively filtering training data (shrivastava2025sample) rather than addressing the root cause of length bias inherent in the basic formulation. We stress that our critique is directed at the specific MDP formulation currently used to model LLM post-training, rather than at RL as a general paradigm for LLMs post-training. This current, restrictive formulation does not fully realize the expressive power of RL. By highlighting these limitations, we aim to encourage the development of richer formulations that can further unlock the potential of RL in LLM post-training.
@@ 2 RELATED WORK
Supervised fine-tuning and reinforcement learning have emerged as the two main paradigms for post-training LLMs. Although early work focused heavily on SFT variants (zelikman2022star; yuan2023scaling; dong2023raft; gulcehre2023reinforced; singh2023beyond), recent efforts have shifted toward RL-inspired methods such as GRPO and its extensions (liu2025understanding; zhang2025grpo; yu2025dapo; yuan2025vapo).
@@
While the original SFT involved finetuning on the traces and solutions supplied externally, a line of work has explored improving language model reasoning via self-generated data, filtering data with verifier and iteratively fine-tuning on the filtered data, a process which we refer to as filtered iterative SFT in our work. Self-Taught Reasoner (STaR) (zelikman2022star) fine-tunes a model on its self generated data by iteratively producing rationales using greedy decoding. It retains only samples whose rationales lead to correct answers. Subsequent works have replaced greedy decoding with temperature sampling as an alternative data generation strategy. Rejection Sampling Fine-Tuning (RFT) (yuan2023scaling) collects reasoning paths by generating multiple outputs from a supervised model and filtering them using binary feedback (accepted/rejected). The model is then fine-tuned once on the accepted samples. RAFT (Reward-ranked Fine-Tuning) (dong2023raft) introduces an iterative framework consisting of three steps: (1) generate a batch of outputs from the model; (2) score the outputs using a reward function and filter high-reward samples and (3) fine-tune the model on the filtered data.
@@
Reinforced Self-Training (ReST) (gulcehre2023reinforced) similarly generates multiple output sequences per prompt, scores them with a reward function, and fine-tunes the model using a reward-weighted loss, rather than hard filtering. ReSTEM (singh2023beyond) extends this setup by fine-tuning the original base model in each iteration, instead of the latest model checkpoint. These methods collectively explore different ways to combine self-generated data, filtering, and reward signals for language model fine-tuning.
@@
The recent and growing work has focused on variations of the GRPO Framework. GRPO (shao2024deepseekmath) introduces group-relative advantages, avoiding state-based value estimation by directly assigning advantages based on relative scores within a group. Several follow-up works build on or modify GRPO in different ways. Dr-GRPO (liu2025understanding) removes standard deviation normalization and sample-length scaling from the GRPO loss, and points out that the objective may introduce a bias toward longer responses. GRPO-Lead (zhang2025grpo) extends GRPO by adding a length-dependent reward to encourage concise outputs, an explicit penalty to reduce low-precision responses, and an advantage reweighting strategy to handle limited rollout data.
In parallel, DAPO (yu2025dapo) introduces a different set of techniques for long chain-of-thought settings. It uses token-level policy gradient loss and includes four key components: Clip-Higher to maintain policy diversity, Dynamic Sampling to improve training stability, Overlong Reward Shaping to reduce noise from long outputs, and the removal of flat-reward samples from training. VAPO (yuan2025vapo) builds on DAPO by introducing value model pretraining, a decoupled and length-adaptive GAE, and group sampling to better handle sparse rewards and variable-length responses.
@@
In this work, We argue that GRPO and its variants, despite their RL framing, rely on structural simplifications such as degenerate MDPs and uniform token-level reward assignment, which effectively reduce the actual learning process to iterative filtered supervised fine-tuning. A very recent work, RAFT++ (xiong2025minimalist) also extended RAFT by incorporating importance sampling and clipping and shows comparable performance as GRPO.
@@
There have also been conflicting claims about the response length: the longer responses have both been celebrated as a sign of improved reasoning guo2025deepseek, and also viewed as a source of inefficiency, with several efforts directed at shortening them. Such efforts include sample-level loss (yu2025dapo; liu2025understanding), length-dependent penalties (yuan2025vapo; aggarwal2025l1; arora2025training), and token-based control, where Thinkless (fang2025thinkless) uses <think>/<short> tokens, while AdaptThink (zhang2025adaptthink), invokes <think> only on difficult questions. However, we show that increase in response length is more likely a side effect of the uniform advantage distribution followed with length scaling. Also, (fatemi2025concise) shows theoretical analysis that generation of longer responses stems from RL-based optimization during training.
@@ 3.2 LLM-MDP and RL for LLMs
Let’s now consider the popular MDP model used for LLMs introduced in DeepSeek-R1 paper (guo2025deepseek), which extends the standard MDP Formulation by incorporating a verifier $V$ into the environment. The resulting formulation is defined as $M_{LLM}=\langle S,A,P,r,\rho,H,V\rangle$. In this scenario, the LLM acts as an agent interacting with an environment whose state space is the set of all possible token sequences up to a maximum length $H$, defined as $S=\bigcup_{k=0}^{H}\mathcal{V}^{k}.$ with $\mathcal{V}$ representing a finite vocabulary of tokens. The action space $A$ is precisely this vocabulary $\mathcal{V}$, corresponding to the selection of individual tokens.
@@
Rewards in this formulation are defined explicitly through an external verifier. Upon episode termination, a reward of 1 is given only if the suffix of the terminal state $s_{H}$, representing the complete sequence generated by the LLM agent, contains the correct solution [^1] In practice, large language models (LLMs) often generate intermediate reasoning traces (e.g., between <Think> tags) followed by final answers (between <Sol> tags), where the <Think> section reflects chain-of-thought reasoning and the <Sol> section is evaluated by a verifier. However, as in the DeepSeek formulation, the state is defined as the sequence of tokens without distinguishing between reasoning and solution parts. We adopt the same simplification and treat the state as the just token sequence, irrespective of internal structure. to the given problem according to the external verifier; otherwise, the reward is zero.
@@
In this setting, fine-tuning an LLM is equivalent to improving the agent’s policy $\pi_{\theta}$ to maximize the expected cumulative reward. RL policy gradient methods are well-suited for this task, as LLMs are transformer-based parameterized neural networks that output a probability distribution over the vocabulary given the context. In the next subsection, we provide an overview of GRPO (shao2024deepseekmath), a policy gradient method used for fine-tuning LLMs.
@@ 3.3 Group Relative Policy Optimization (GRPO)
$\mathcal{J}(\theta)$ $=\mathbb{E}_{q\sim P(Q),\,\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O|q)}$ (1)
$\quad\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\Biggl\{\min\Biggl(\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}\hat{A}_{i,t},\text{clip}\left(\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})},1-\varepsilon,1+\varepsilon\right)\hat{A}_{i,t}\Biggr)$
$\quad-\beta\mathbb{D}_{KL}\left[\pi_{\theta}\|\pi_{\text{ref}}\right]\Biggr\}$
@@ 4.1 Simplifying the GRPO Objective
$\mathcal{J}(\theta)$ $=\mathbb{E}_{q\sim P(Q),\,\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O|q)}$ (2)
$\quad\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\Biggl\{\min\Biggl(\mathcal{ISR}_{i,t}(\theta)\hat{A}_{i,t},\text{clip}\Biggl(\mathcal{ISR}_{i,t}(\theta),1-\varepsilon,1+\varepsilon\Biggr)\hat{A}_{i,t}\Biggr)-\beta\mathbb{D}_{KL}\left[\pi_{\theta}\|\pi_{\text{ref}}\right]\Biggr\}$
@@
Recent works ((liu2025understanding), (yu2025dapo), (yuan2025vapo)) show that the KL penalty term has limited effect on performance, since the clipping operation already keeps the updated policy within the trust region of the old policy ((schulman2015trust), (schulman2017proximal)). Therefore, we can relax the KL penalty term from the objective, simplifying it to:

### v4

# RL in Name Only? Analyzing the Structural Assumptions in RL post-training for LLMs

@@ Abstract
Reinforcement learning based post-training of large language models (LLMs) has recently gained attention, particularly following the release of DeepSeek R1 (Guo et al., 2025), which applied GRPO for fine-tuning. Amid the growing claims around improved reasoning abilities attributed to RL post-training, we critically examine the formulation and assumptions underlying these methods. We start by highlighting popular structural assumptions made in modeling LLM training as an MDP, and show how they lead to a degenerate MDP, that characterizes the problem as a contextual bandit, where RL updates naturally collapse into a form of on-policy variant of outcome-driven supervised learning. The two critical structural assumptions include (1) making the MDP states be just a concatenation of the actions with states becoming the context window and the actions becoming the tokens in LLMs and (2) splitting the reward of a state-action trajectory uniformly across the trajectory. Our comprehensive analysis demonstrates that, due to these simplifying assumptions, GRPO objective reduces to filtered Iterative SFT, an on-policy variant of supervised fine-tuning. Our experiments on benchmarks including GSM8K and Countdown, across a diverse set of model families show that Filtered Iterative SFT, incorporating both positive and negative samples, achieves performance comparable to GRPO-based training. We also show that these structural assumptions indirectly incentivize RL to generate longer sequences of intermediate tokens which in turn feeds into the narrative of “RL incentivizing thinking because it generates longer thinking traces.”
@@ 1 Introduction
In this work, we critically examine the foundations of applying RL to LLM post-training through the lens of the specific MDP formulation that has been popularly used to model LLMs  (Guo et al., 2025). Under this formulation, each state is a sequence of previously generated tokens, transitions are deterministic concatenations of tokens, rewards are assigned only at the terminal state based on external verification of correctness, and the credit assignment problem is short-circuited by effectively dividing this terminal reward across all tokens (or actions) equally. These structural assumptions sharply distinguish the setting from the general MDP framework typically considered in RL, characterizing it instead as a contextual bandit problem. In this view, the input prompt serves as the context, while the generation of a complete sequence, comprising both intermediate and solution tokens, represents the pulling of a single arm from a vast space of possible responses. The quality of the selected arm is defined solely by an external verifier based on the solution’s correctness. Consequently, token generation acts as the constructive formation of an arm rather than a navigation through an environment with intermediate state changes.
@@
Our contributions are twofold. Firstly, we theoretically show that under these structural assumptions, the GRPO objective naturally collapses into an on-policy variant of iterative filtered SFT. By decoupling and simplifying the GRPO objective for trajectories labeled as positive (correct solutions) and negative (incorrect solutions), we demonstrate that on-policy variant of iterative supervised fine-tuning (SFT) guided by an external verifier yields comparable performance. This equivalence highlights that the core value of RL in this regime lies in on-policy exploration and outcome-based filtering rather than temporal credit assignment. The method effectively performs a search over the base model’s solution space, which is then distilled into the policy via SFT-style updates. This further implies that current RL methods may not be imparting genuinely novel reasoning capabilities, but are instead strategically leveraging and refining capabilities already inherent within the pretrained models Yue et al. (2025).
@@
Secondly, we empirically demonstrate that the terminal reward distribution assumption, wherein the relative advantage score is uniformly allocated across all tokens in a response, directly contributes to the increasing response lengths observed throughout training. Contrary to the claims made in the DeepSeek-R1 (Guo et al., 2025) paper, which attributes the increased response length at test-time to computational scaling, self-reflection, self-verification, and back-tracking, we show that the primary driver of response elongation is this uniform credit distribution. Since each token is incentivized equally without considering the "return-to-go" from any intermediate state, RL-based methods like GRPO inadvertently encourage verbose outputs.
@@
This paper thus aims to clarify the misconceptions regarding the role of RL in enhancing LLM capabilities by providing a comprehensive theoretical and empirical analysis showing that idiosyncratic assumptions made in formulating the MDP influence model behavior during post training. A key implication of our analysis is that misunderstanding the consequences of these structural assumptions has led to research directions that attempt to fix superficial symptoms such as introducing length penalties (Arora and Zanette, 2025) or selectively filtering training data (Shrivastava et al., 2025) whereas we highlight the underlying root cause inherent in the basic formulation. Under the equivalent FISFT formulation, the resulting updates do not depend on advantage estimates derived from reward signals, and FISFT variants, using a verifier signal alone can achieve comparable performance. By making the GRPO-FISFT equivalence explicit, we provide a theoretical foundation for recent empirical findings like increase in response length, and the underlying training dynamics from the lens of FISFT.
@@ 2 Related Work
Supervised fine-tuning and reinforcement learning have emerged as the two main paradigms for post-training LLMs. Although early work focused heavily on SFT variants (Zelikman et al., 2022; Yuan et al., 2023; Dong et al., 2023; Gulcehre et al., 2023; Singh et al., 2023), recent efforts have shifted toward RL-inspired methods such as GRPO and its extensions (Liu et al., 2025; Zhang and Zuo, 2025; Yu et al., 2025; Yuan et al., 2025).
@@
While the original SFT involved finetuning on the traces and solutions supplied externally, a line of work has explored improving language model reasoning via self-generated data, filtering data with verifier and iteratively fine-tuning on the filtered data, a process which we refer to as filtered iterative SFT in our work. Self-Taught Reasoner (STaR) (Zelikman et al., 2022) fine-tunes a model on its self generated data by iteratively producing rationales using greedy decoding. It retains only samples whose rationales lead to correct answers. Subsequent works have replaced greedy decoding with temperature sampling as an alternative data generation strategy. Rejection Sampling Fine-Tuning (RFT) (Yuan et al., 2023) collects reasoning paths by generating multiple outputs from a supervised model and filtering them using binary feedback (accepted/rejected). The model is then fine-tuned once on the accepted samples. RAFT (Reward-ranked Fine-Tuning) (Dong et al., 2023) introduces an iterative framework consisting of three steps: (1) generate a batch of outputs from the model; (2) score the outputs using a reward function and filter high-reward samples and (3) fine-tune the model on the filtered data.
@@
Reinforced Self-Training (ReST) (Gulcehre et al., 2023) similarly generates multiple output sequences per prompt, scores them with a reward function, and fine-tunes the model using a reward-weighted loss, rather than hard filtering. ReSTEM (Singh et al., 2023) extends this setup by fine-tuning the original base model in each iteration, instead of the latest model checkpoint. These methods collectively explore different ways to combine self-generated data, filtering, and reward signals for language model fine-tuning.
@@
GRPO (Shao et al., 2024) introduces group-relative advantages, avoiding state-based value estimation by directly assigning advantages based on relative scores within a group. Several follow-up works build on or modify GRPO in different ways. Dr-GRPO (Liu et al., 2025) removes standard deviation normalization and sample-length scaling from the GRPO loss, and points out that the objective may introduce a bias toward longer responses. GRPO-Lead (Zhang and Zuo, 2025) extends GRPO by adding a length-dependent reward, and an advantage reweighting strategy.In parallel, DAPO (Yu et al., 2025) introduces a different set of techniques for long chain-of-thought settings. VAPO (Yuan et al., 2025) builds on DAPO by introducing value model pretraining, a decoupled and length-adaptive GAE, and group sampling.
@@
In this work, we argue that GRPO and its variants, despite their RL framing, rely on structural simplifications such as degenerate MDPs and uniform token-level reward assignment, which effectively reduce the actual learning process to iterative filtered supervised fine-tuning. A very recent work, RAFT++ (Xiong et al., 2025) also extended RAFT by incorporating importance sampling and clipping, and provides ablations on different components of GRPO.
Further, Thought Anchors Bogdan et al. (2025) conducted counterfactual analyses not all tokens contribute equally to producing the correct answer. Their work identifies pivotal tokens/sentences that are particularly influential in reaching the correct final answer using attention patterns. In contrast, GRPO assigns equal weight to all tokens,
@@
There have also been conflicting claims about the response length: the longer responses have both been celebrated as a sign of improved reasoning Guo et al. (2025), and also viewed as a source of inefficiency, with several efforts directed at shortening them. Such efforts include sample-level loss (Yu et al., 2025; Liu et al., 2025), length-dependent penalties (Yuan et al., 2025; Aggarwal and Welleck, 2025; Arora and Zanette, 2025), and token-based control, where Thinkless (Fang et al., 2025) uses <think>/<short> tokens, while AdaptThink (Zhang et al., 2025), invokes <think> only on difficult questions. However, we show that increase in response length is more likely a side effect of the uniform advantage distribution followed with length scaling. Also, (Fatemi et al., 2025) shows theoretical analysis that generation of longer responses stems from RL-based optimization during training.
@@ 3.2 LLM-MDP and RL for LLMs
Let’s now consider the popular MDP model used for LLMs introduced in DeepSeek-R1 paper (Guo et al., 2025), which extends the standard MDP Formulation by incorporating a verifier $V$ into the environment. The resulting formulation is defined as $M_{LLM}=\langle S,A,P,r,\rho,H,V\rangle$. In this scenario, the LLM acts as an agent interacting with an environment whose state space is the set of all possible token sequences up to a maximum length $H$, defined as $S=\bigcup_{k=0}^{H}\mathcal{V}^{k}.$ with $\mathcal{V}$ representing a finite vocabulary of tokens. The action space $A$ is precisely this vocabulary $\mathcal{V}$, corresponding to the selection of individual tokens.
@@
Rewards in this formulation are defined explicitly through an external verifier. Upon episode termination, a reward of 1 is given only if the suffix of the terminal state $s_{H}$, representing the complete sequence generated by the LLM agent, contains the correct solution [^1] In practice, large language models (LLMs) often generate intermediate reasoning traces (e.g., between <Think> tags) followed by final answers (between <Sol> tags), where the <Think> section reflects chain-of-thought reasoning and the <Sol> section is evaluated by a verifier. However, as in the DeepSeek formulation, the state is defined as the sequence of tokens without distinguishing between reasoning and solution parts. We adopt the same simplification and treat the state as the just token sequence, irrespective of internal structure. to the given problem according to the external verifier; otherwise, the reward is 0.
@@
Under this LLM-MDP formulation, RL-based post-training methods frame fine-tuning as an optimization problem where the goal is to improve the agent’s policy $\pi_{\theta}$, to maximize expected cumulative rewards. Because LLMs are transformer-based parameterized neural networks that output a probability distribution over a vocabulary/action space given a specific context, these methods typically employ policy gradient algorithms, such as PPO and its variants, for policy improvement.
We argue that the structural assumptions inherent in this LLM-MDP formulation effectively collapse the sequential decision-making process into a contextual bandit setting. In this view, the initial state $s_{0}$ (the question prompt) serves as the context in the bandit formulation. Furthermore, because transitions are deterministic and each state $s_{t}$ encodes the full history of preceding tokens, the generation of a complete sequence $s_{H}$ can be viewed as pulling a single arm or selecting a macro-action from the vast space of possible responses $\mathcal{V}^{H}$. The terminal reward $r(s_{H})$ then acts as bandit feedback observed only after the full macro-action is complete.
Consequently, the sequential nature of token generation functions merely as a mechanism to construct this arm, rather than as a navigation through a stochastic environment with intermediate state changes. We contend that because these structural assumptions cause the MDP to behave as a contextual bandit problem, the GRPO objective naturally becomes equivalent to an on-policy variant of Filtered-SFT, that we theoretically prove in section 4. In the next subsection, we first provide an overview of GRPO (Shao et al., 2024), a policy gradient method used for fine-tuning LLMs.
@@ 3.3 Group Relative Policy Optimization (GRPO)
$\mathcal{J}(\theta)$ $=\mathbb{E}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\Biggl\{\min\Biggl(\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})}\hat{A}_{i,t},\text{clip}\left(\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}|q,o_{i,<t})},1-\varepsilon,1+\varepsilon\right)\hat{A}_{i,t}\Biggr)$ (1)
$\quad-\beta\mathbb{D}_{\mathrm{KL}}\left[\pi_{\theta}\|\pi_{\text{ref}}\right]\Biggr\}\Bigg]$
@@ 4.1 Simplifying the GRPO Objective
$\mathcal{J}(\theta)$ $=\mathbb{E}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\Biggl\{\min\Biggl(\mathcal{ISR}_{i,t}(\theta)\hat{A}_{i,t},\text{clip}\Biggl(\mathcal{ISR}_{i,t}(\theta),1-\varepsilon,1+\varepsilon\Biggr)\hat{A}_{i,t}\Biggr)$ (2)
$\quad-\beta\mathbb{D}_{KL}\left[\pi_{\theta}\|\pi_{\text{ref}}\right]\Biggr\}\Bigg]$
@@
Recent works ((Liu et al., 2025), (Yu et al., 2025), (Yuan et al., 2025)) show that the KL penalty term has limited effect on performance, since the clipping operation already keeps the updated policy within the trust region of the old policy ((Schulman et al., 2015), (Schulman et al., 2017)). Therefore, we can relax the KL penalty term from the objective, simplifying it to: